# Reflectance Prediction-based Knowledge Distillation for Robust 3D Object Detection in Compressed Point Clouds

### v1

@@ III-C Reflectance Prediction Module
Concerning the reflectance reconstruction of compressed point clouds, Unicorn [7] utilizes current-scale reflectance residual features and low-scale reflectance priors through a DNN module to predict reflectance. In contrast, our RP module exploits the geometric features of non-reflectance compressed point clouds and the reflectance priors of raw point clouds to achieve precise reflectance reconstruction at the receiver. This innovative design not only reduces the transmission rate in compression transmission systems but also enhances the robustness of 3D detectors against non-reflectance compressed point clouds. As shown in Fig. 3, the structure and principle of our RP module are detailed as follows.
@@
where denotes a sparse tensor transformation module, is the 3D sparse convolution with a downsampling size (i.e., the initial size) in the voxel-wise basic backbone, and represents the voxel set extraction corresponding to . The voxel set extraction structure in our RP module is identical to that in PV-RCNN [22], but the number of sampling points is equal to the number of compressed points in the current frame. Since the reflectance is a single-channel floating-point scalar in the range of , we use two fully connected (FC) layers as a prediction head to generate compressed point reflectance, denoted as . The predicted reflectance of compressed points, , is then incorporated into the corresponding compressed point-cloud voxels based on their spatial relationship for subsequent training and detection.
We screen the compressed point reflectance labels with values greater than zero, and define their indices as . We apply Mean Squared Error Loss (MSELoss) to constrain compressed point reflectance prediction, and the corresponding loss function is expressed as follows:
where represents the number of compressed point reflectance labels with values greater than zero.
@@
Subsequently, we predict raw point reflectance and incorporate it into corresponding voxels to obtain raw point-cloud voxels with the recovered reflectance .
@@
where represents the number of raw point reflectance labels with values greater than zero.
III-D Reflectance Prediction-based Knowledge Distillation
In lossy compression, non-reflectance compressed point clouds undergo point number reduction and reconstruction errors, leading to incomplete and biased geometric information. If only these compressed point clouds are used for reflectance prediction and object detection, it is difficult to match the detection accuracy of raw point clouds. Drawing from the weak-data robustness research [53, 54, 55] in 3D object detection knowledge distillation, we treat the non-reflectance compressed point clouds as weak students, enabling them to learn accurate and complete knowledge from raw point clouds to improve detection performance for compressed point clouds. Specifically, we propose two components: RKD, which transfers reflectance prediction knowledge from our RP module, and DKD, which refines the first-stage proposals in the two-stage 3D object detector, as depicted in Fig. 3. Both proposed distillation methods fall under the response-based knowledge distillation category, and their principles are outlined as follows.
Reflectance Knowledge Distillation. Building upon the RP module’s prediction outputs, we obtain the compressed point reflectance and raw point reflectance . Since they do not have a one-to-one correspondence, we map the recovered teacher reflectance from raw point-cloud voxels to corresponding compressed points. The mapped teacher reflectance of each compressed point serves as a soft label of RKD. According to the spatial relationship in the RCM module, the raw point-cloud voxel corresponding to compressed point coordinates is denoted as , and the RKD soft labels are obtained by the following formula:
where represents the mapped reflectance value from a raw point-cloud voxel to corresponding compressed point coordinates , and is the number of compressed points that can be successfully matched with raw point-cloud voxels in the RCM module. We also adopt MSELoss as the loss function to constrain the reflectance distillation process, and the RKD loss function is formulated as
where denotes screening foreground compression points during loss computation, and is the number of foreground compression points.
Detection Knowledge Distillation. In Fig. 3, the two-stage 3D detector transfers key features from the first-stage network to the second-stage detector for detection refinement. Additionally, the first-stage dense proposals passed to the second stage inherently carry valuable detection knowledge for distillation. DKD combines the compressed point-cloud proposals and raw point-cloud proposals to form a student-teacher pair for response-based knowledge distillation. In these proposals, and represent the sigmoid-activated classification scores and bounding boxes, respectively. For distillation constraints, we apply Logit KD proposed in SparseKD [45] to design the distillation loss. Therefore, the DKD loss function can be expressed as:
@@ III-E Loss Function
In this study, we pre-train separate detectors for the compressed and raw point clouds and initialize the training parameters of our RPKD framework using these pre-trained models. Therefore, the compressed point-cloud detection loss and the raw point-cloud detection loss are defined as follows:
@@ IV Experiments
For 3D object detection in compressed point clouds, we utilized LiDAR point clouds from the KITTI dataset and Waymo Open Dataset (WOD) as raw point-cloud inputs for the compression network. Subsequently, we compressed and reconstructed the geometric coordinates of these point clouds at multiple code rates through deep learning-based PCC-S, resulting in non-reflectance compressed point clouds. The proposed RPKD framework was trained on both compressed and raw point clouds, using only compressed point clouds for inference. Extensive experiments demonstrate that our method significantly enhances detection robustness for non-reflectance compressed point clouds generated by voxel-based compression. The following sections provide a comprehensive explanation of the implementation details, result comparisons, and ablation analysis.
| Data Type | Method | Car 3D | Ped. 3D | Cyc. 3D | mAP | ||||||
| Easy | Mod. | Hard | Easy | Mod. | Hard | Easy | Mod. | Hard | |||
| Raw Data | PV-RCNN | 91.97 | 84.72 | 82.32 | 68.60 | 60.61 | 55.55 | 92.06 | 74.48 | 69.63 | 75.56 |
| RP-PV(T2) | 92.10 | 84.36 | 82.29 | 63.62 | 56.35 | 50.88 | 90.30 | 70.75 | 66.96 | 73.07 | |
| Voxel-RCNN | 92.55 | 84.65 | 82.41 | 66.07 | 59.30 | 53.73 | 92.81 | 73.55 | 68.97 | 74.89 | |
| RP-Voxel(T3) | 92.18 | 84.12 | 81.75 | 62.04 | 56.11 | 51.05 | 93.67 | 72.87 | 68.41 | 73.58 | |
| PCC-S-C12 (Bpp=3.864) | PV-RCNN(Baseline) | 91.48 | 82.51 | 80.72 | 56.17 | 49.66 | 45.37 | 87.85 | 68.51 | 64.21 | 69.61 |
| RPKD-PV(Ours) | 92.02 | 83.13 | 80.21 | 65.72 | 58.68 | 53.72 | 90.39 | 68.63 | 63.92 | 72.94 | |
| Voxel-RCNN(Baseline) | 92.27 | 82.79 | 80.17 | 63.64 | 56.21 | 51.21 | 90.69 | 71.35 | 66.92 | 72.81 | |
| RPKD-Voxel(Ours) | 92.06 | 82.91 | 79.98 | 70.40 | 64.14 | 57.63 | 92.26 | 70.31 | 65.84 | 75.06 | |
| PCC-S-C11 (Bpp=2.146) | PV-RCNN(Baseline) | 90.85 | 82.25 | 80.33 | 57.79 | 50.83 | 46.20 | 88.92 | 65.79 | 61.40 | 69.37 |
| RPKD-PV(Ours) | 91.38 | 82.95 | 79.96 | 70.93 | 62.81 | 56.39 | 89.21 | 66.68 | 62.09 | 73.60 | |
| Voxel-RCNN(Baseline) | 92.03 | 82.72 | 80.16 | 61.49 | 54.77 | 49.30 | 88.27 | 67.99 | 63.97 | 71.19 | |
| RPKD-Voxel(Ours) | 91.78 | 83.00 | 80.30 | 70.33 | 63.36 | 57.17 | 92.12 | 68.10 | 63.54 | 74.41 | |
| PCC-S-C10 (Bpp=1.047) | PV-RCNN(Baseline) | 90.74 | 81.40 | 79.24 | 54.71 | 47.58 | 44.12 | 81.08 | 61.64 | 57.45 | 66.44 |
| RPKD-PV(Ours) | 91.53 | 82.52 | 79.45 | 63.24 | 57.24 | 51.90 | 85.51 | 61.70 | 57.68 | 70.09 | |
| Voxel-RCNN(Baseline) | 90.96 | 81.37 | 78.73 | 60.10 | 51.66 | 45.88 | 83.05 | 62.86 | 58.62 | 68.14 | |
| RPKD-Voxel(Ours) | 90.67 | 81.87 | 78.95 | 65.95 | 59.02 | 53.54 | 87.79 | 64.80 | 60.35 | 71.44 | |
@@ IV-A1 Datasets and Hardware
For the KITTI dataset, we used 7481 labeled training samples as raw point clouds, dividing them into a training split with 3712 samples and a validation split with 3769 samples. Firstly, the PCC-S network employed the training split of raw point clouds for compression learning and reconstructed 7481 frames of point clouds through the trained compression model at octree levels of 12, 11, and 10, respectively. Regarding the training configuration, we trained the PCC-S network for 20 epochs on a single RTX 8000 GPU with a batch size of 2 and a maximum octree level of 12. Then, we established independent models for each code rate and conducted joint training on both the compressed and raw point clouds to build robust 3D detectors for compressed point clouds. These detectors were trained for 80 epochs on 4 RTX 4090 GPUs with a batch size of 16.
WOD is a large-scale detection dataset composed of 798 training sequences and 202 validation sequences. Due to the limited training efficiency of PCC-S and the computing power of a single RTX 8000 GPU, we only selected 20 training sequences and 10 validation sequences for research. Thus, the raw point clouds consisted of a training split with 3966 samples and a validation split with 1967 samples. Using the same training configuration as KITTI, we obtained 5933 frames of compressed point clouds by PCC-S on WOD at octree levels 12, 11, and 10, respectively. Additionally, the detectors of our RPKD framework were trained for 30 epochs on 4 RTX 4090 GPUs with a batch size of 8.
@@ IV-A2 Training Parameters
In the RCM module, we matched compressed points with raw point-cloud voxels within a 1-voxel matching range, corresponding to spatial voxels. During pre-training, the compressed and raw point-cloud detectors shared the same loss weights, and these loss weights were set to . In the knowledge distillation stage, we designed two distillation methods, RKD and DKD, whose loss weights were set to . Specifically, the distillation constraints for DKD consisted of classification distillation loss and bounding box distillation loss , and their loss weights were . For dataset configuration, we trained compressed point-cloud detectors using the standard parameters of public datasets and expanded single-source data to two-source data for joint processing.
@@ IV-A3 Evaluation Metrics
We evaluated the robustness of our RPKD framework for 3D object detection on compressed point clouds using the KITTI and WOD validation splits. The primary metric for KITTI was average precision (AP) at 40 recall positions with 3D IoU thresholds of 0.7, 0.5, and 0.5 for cars, pedestrians, and cyclists, respectively. For WOD, the metrics were AP and average precision weighted by heading (APH). To compare performance across multiple code rates and methods, we used mean Average Precision (mAP) and mean APH (mAPH) across categories and difficulty levels. The Bit-per-point (Bpp) metric represents the average number of bits required to encode each point, indicating the code rate of compressed point clouds.
IV-A4 Controlled Experiment Design
Firstly, in terms of data types, raw data, PCC-S-C12, PCC-S-C11, and PCC-S-C10 represented raw point clouds, the compressed point clouds of octree levels 12, 11, and 10, respectively. Then, the detection performance of PV-RCNN at each code rate served as the baseline, and the designed control experiments involved three detection methods, namely SparseKD-PV, RPV-KD, and RPKD-PV. In SparseKD-PV, with the teacher detector of PV-RCNN (T1) on raw data, we utilized Logit KD and Label KD for distillation learning and performance evaluation of compressed point-cloud detectors. RPV-KD and SparseKD-PV share the same teacher model and distillation methods, while the key difference is that RPV-KD mapped raw point reflectance into corresponding compressed points based on the matching principle of our RCM module (Section III-B) to simulate compressed point clouds with decoded reflectance. RPKD-PV denotes the implementation of our method on the PV-RCNN backbone, with RP-PV (T2) as the teacher detector. Finally, referring to the experimental settings of RPKD-PV, we performed generalization verification of our method on the Voxel-RCNN backbone with RPKD-Voxel, using RP-Voxel (T3) as the teacher detector.
| Data Type | Detection Method | Vehicle(L1) | Vehicle(L2) | Pedestrian(L1) | Pedestrian(L2) | mAP/mAPH |
| Raw Data | PV-RCNN | 61.37/60.01 | 53.20/52.00 | 65.69/33.70 | 59.00/30.27 | 59.82/43.99 |
| RP-PV(T2) | 61.75/60.76 | 53.44/52.56 | 60.42/29.80 | 54.14/26.70 | 57.44/42.46 | |
| PCC-S-C12 (Bpp=4.637) | PV-RCNN(Baseline) | 61.27/60.07 | 52.87/51.81 | 59.89/30.98 | 53.60/27.78 | 56.91/42.66 |
| RPKD-PV(Ours) | 62.08/61.02 | 53.72/52.77 | 59.62/34.44 | 53.37/30.82 | 57.20/44.76 | |
| PCC-S-C11 (Bpp=2.711) | PV-RCNN(Baseline) | 60.51/59.52 | 52.30/51.43 | 58.75/29.72 | 52.57/26.61 | 56.03/41.82 |
| RPKD-PV(Ours) | 61.41/60.38 | 53.22/52.30 | 58.33/32.55 | 52.21/29.14 | 56.29/43.59 | |
| PCC-S-C10 (Bpp=1.359) | PV-RCNN(Baseline) | 60.09/58.81 | 51.95/50.82 | 56.76/26.30 | 50.77/23.53 | 54.89/39.86 |
| RPKD-PV(Ours) | 61.13/60.04 | 52.92/51.95 | 57.49/28.16 | 51.45/25.20 | 55.75/41.34 |

### v2

@@ III-C Reflectance Prediction Module
Concerning the reflectance reconstruction of compressed point clouds, Unicorn [7] utilized current-scale reflectance residual features and low-scale reflectance priors through a DNN module to predict reflectance. In contrast, our RP module exploits the geometric features of non-reflectance compressed point clouds and the reflectance priors of raw point clouds to reconstruct reflectance at the receiver. In collaborative perception with LiDAR point cloud compression, the concept of reflectance prediction is crucial for ensuring detection robustness against low-quality compressed data, while eliminating reflectance encoding to push the boundaries of transmission bandwidth reduction. As shown in Fig. 3, the structure and principle of our RP module are detailed as follows.
@@
where denotes a sparse tensor transformation module, is the 3D sparse convolution with a downsampling size (i.e., the initial size) in the voxel-wise basic backbone, and represents the voxel set extraction corresponding to . The voxel set extraction structure in our RP module is identical to that in PV-RCNN [23], but the number of sampling points is equal to the number of compressed points in the current frame. Since the reflectance input to the detection network is a single-channel floating-point scalar in the range of , we use two fully connected (FC) layers as a prediction head to generate compressed point reflectance. The predicted reflectance is then incorporated into the corresponding compressed point-cloud voxels based on their spatial relationship for subsequent training and detection.
We screen the compressed point reflectance labels with values greater than zero, and define their indices as . Mean Squared Error Loss (MSELoss) is applied to constrain compressed point reflectance prediction, and the corresponding loss function is expressed as follows:
where represents the number of compressed point reflectance labels with values greater than zero, and denotes the predicted reflectance for compressed points.
@@
where represents raw point coordinates, and is the number of raw points. Subsequently, we predict raw point reflectance and assign it to corresponding voxels, resulting in raw point-cloud voxels with recovered reflectance.
@@
where represents the number of raw point reflectance labels with values greater than zero, and denotes the predicted reflectance for raw points.
III-D Reflectance Prediction-based CDTS
Inspired by weak-data robustness research [54, 15, 55], the proposed CDTS employs raw and compressed data models in a teacher–student framework for knowledge distillation (as shown in Fig. 5), thereby improving detection performance on low-quality compressed data. To tackle the critical challenge of missing LiDAR reflectance in lossy geometric compression, our RPKD (CDTS) strengthens reflectance reconstruction through the integration of reflectance priors, the RP module, and the RKD constraint. In knowledge distillation, we introduce two components: RKD, which transfers reflectance prediction knowledge from the RP module, and DKD, which refines first-stage proposals in the two-stage 3D object detector. Both methods belong to response-based knowledge distillation, and their principles are outlined below.
Reflectance Knowledge Distillation. Building upon the RP module’s outputs, we obtain the predicted raw and compressed point reflectance. Since they do not have a one-to-one correspondence, we map the recovered teacher reflectance from raw point-cloud voxels to corresponding compressed points. The mapped teacher reflectance of each compressed point serves as a soft label of RKD. We also adopt MSELoss to constrain the reflectance distillation process, and the RKD loss function is formulated as
where denotes screening foreground compression points during loss computation, is the number of compressed points that can be successfully matched with raw point-cloud voxels, and is the number of foreground compression points.
Detection Knowledge Distillation. In Fig. 3, the two-stage 3D detector transfers key features from the first-stage network to the second-stage detector for detection refinement. Additionally, the first-stage dense proposals passed to the second stage inherently carry valuable detection knowledge for distillation. DKD combines the raw and compressed point-cloud proposals to form a teacher-student pair for response-based knowledge distillation. For distillation constraints, we apply Logit KD proposed in SparseKD [16] to design the distillation loss. Therefore, the DKD loss function can be expressed as:
@@ III-E Loss Function
In this study, we pretrain separate detectors on compressed and raw point clouds and initialize the RPKD framework parameters with these pretrained models. Their loss functions are defined as follows:
@@ IV Experiments
For 3D object detection in compressed point clouds, we employed LiDAR data from the KITTI and DAIR-V2X-V datasets as raw point-cloud inputs for the compression network. The geometric coordinates of these point clouds were compressed and reconstructed at multiple code rates using the voxel-based lossy geometric compression method PCC-S [9]. Building on the proposed CDTS, extensive experiments demonstrate that our RPKD (CDTS) significantly enhances detection robustness for non-reflectance compressed point clouds. The following sections detail the implementation, comparative results, and ablation analysis.
| Data Type | Bpp | Method | Car 3D | Ped. 3D | Cyc. 3D | mAP | ||||||
| Easy | Mod. | Hard | Easy | Mod. | Hard | Easy | Mod. | Hard | ||||
| Raw Data | - | PV-RCNN [23]* | 91.97 | 84.72 | 82.32 | 68.60 | 60.61 | 55.55 | 92.06 | 74.48 | 69.63 | 75.55 |
| PCC-S-C12 | 3.86 | PV-RCNN (STS-C) | 91.48 | 82.51 | 80.72 | 56.17 | 49.66 | 45.37 | 87.85 | 68.51 | 64.21 | 69.61 |
| RPKD-PV (CDTS) | 92.02 | 83.13 | 80.21 | 65.72 | 58.68 | 53.72 | 90.39 | 68.63 | 63.92 | 72.94 | ||
| PCC-S-C11 | 2.15 | PV-RCNN (STS-C) | 90.85 | 82.25 | 80.33 | 57.79 | 50.83 | 46.20 | 88.92 | 65.79 | 61.40 | 69.37 |
| RPKD-PV (CDTS) | 91.38 | 82.95 | 79.96 | 70.93 | 62.81 | 56.39 | 89.21 | 66.68 | 62.09 | 73.60 | ||
| PCC-S-C10 | 1.05 | PV-RCNN (STS-C) | 90.74 | 81.40 | 79.24 | 54.71 | 47.58 | 44.12 | 81.08 | 61.64 | 57.45 | 66.44 |
| RPKD-PV (CDTS) | 91.53 | 82.52 | 79.45 | 63.24 | 57.24 | 51.90 | 85.51 | 61.70 | 57.68 | 70.09 | ||
| Raw Data | - | Voxel-RCNN [24]* | 92.55 | 84.65 | 82.41 | 66.07 | 59.30 | 53.73 | 92.81 | 73.55 | 68.97 | 74.89 |
| PCC-S-C12 | 3.86 | Voxel-RCNN (STS-C) | 92.27 | 82.79 | 80.17 | 63.64 | 56.21 | 51.21 | 90.69 | 71.35 | 66.92 | 72.81 |
| RPKD-Voxel (CDTS) | 91.89 | 82.94 | 80.03 | 71.49 | 64.33 | 58.50 | 92.28 | 71.47 | 66.85 | 75.53 | ||
| PCC-S-C11 | 2.15 | Voxel-RCNN (STS-C) | 92.03 | 82.72 | 80.16 | 61.49 | 54.77 | 49.30 | 88.27 | 67.99 | 63.97 | 71.19 |
| RPKD-Voxel (CDTS) | 91.74 | 82.90 | 80.22 | 69.65 | 62.83 | 56.10 | 91.80 | 68.36 | 63.94 | 74.17 | ||
| PCC-S-C10 | 1.05 | Voxel-RCNN (STS-C) | 90.96 | 81.37 | 78.73 | 60.10 | 51.66 | 45.88 | 83.05 | 62.86 | 58.62 | 68.14 |
| RPKD-Voxel (CDTS) | 91.37 | 82.33 | 79.50 | 66.72 | 59.53 | 53.88 | 84.68 | 63.03 | 58.93 | 71.11 | ||
| Raw Data | - | SECOND [25]* | 90.43 | 81.50 | 78.50 | 55.89 | 50.97 | 46.38 | 81.91 | 64.40 | 60.53 | 67.83 |
| PCC-S-C12 | 3.86 | SECOND (STS-C) | 90.34 | 81.21 | 78.32 | 55.59 | 49.67 | 45.32 | 80.19 | 65.19 | 61.23 | 67.45 |
| RPKD-SECOND (CDTS) | 91.08 | 81.96 | 78.87 | 58.42 | 53.06 | 48.55 | 86.00 | 68.38 | 64.35 | 70.07 | ||
| PCC-S-C11 | 2.15 | SECOND (STS-C) | 90.17 | 80.90 | 78.09 | 52.58 | 47.66 | 43.16 | 81.92 | 61.21 | 57.31 | 65.89 |
| RPKD-SECOND (CDTS) | 90.87 | 81.69 | 78.62 | 57.35 | 52.88 | 48.44 | 84.84 | 66.07 | 61.79 | 69.17 | ||
| PCC-S-C10 | 1.05 | SECOND (STS-C) | 88.25 | 78.80 | 75.72 | 50.48 | 45.29 | 41.37 | 75.59 | 56.28 | 52.66 | 62.72 |
| RPKD-SECOND (CDTS) | 90.08 | 80.74 | 77.52 | 56.76 | 52.34 | 47.71 | 81.81 | 62.39 | 58.32 | 67.52 | ||
@@ IV-A1 Datasets and Hardware
In KITTI, we selected 7,481 labeled training samples as raw point clouds, dividing them into a training split of 3,712 samples and a validation split of 3,769 samples. Similarly, in DAIR-V2X-V, we selected 6,509 training frames, which were split into 4,335 frames for training and 2,174 frames for validation. We trained the PCC-S network on the raw point-cloud training split for compression learning and applied to reconstruct point clouds at octree levels 12, 11, and 10, denoted as PCC-S-C12, PCC-S-C11, and PCC-S-C10, respectively. The compression network was optimized for 20 epochs on a single RTX 8000 GPU with a batch size of 2 and a maximum octree level of 12. Independent detection models for each code rate were trained for 80 epochs on 4 RTX 4090 GPUs with a batch size of 16.
@@ IV-A2 Training Parameters
In the RCM module, we matched compressed points with raw point-cloud voxels within a 1-voxel matching range, corresponding to spatial voxels. During pretraining, the raw and compressed point-cloud detectors shared the same loss weights, and these loss weights were set to . For CDTS, we introduced two distillation methods, namely RKD and DKD, whose loss weights were configured as . Specifically, the distillation constraints for DKD included classification distillation loss and bounding box distillation loss , and their loss weights were .
@@ IV-A3 Evaluation Metrics
We evaluated the robustness of our RPKD framework for 3D object detection on compressed point clouds using the KITTI and DAIR-V2X-V validation splits. For KITTI and DAIR-V2X-V, the primary metric is average precision (AP) at 40 recall positions with 3D IoU thresholds of 0.7, 0.5, and 0.5 for car, pedestrian, and cyclist categories, respectively. The Bit-per-point (Bpp) metric represents the average number of bits required to encode each point, indicating the code rate of compressed point clouds.
| Data Type | Bpp | Method | Car 3D | Ped. 3D | Cyc. 3D | mAP | ||||||
| Easy | Mod. | Hard | Easy | Mod. | Hard | Easy | Mod. | Hard | ||||
| Raw Data | - | PV-RCNN [23]* | 72.06 | 60.50 | 57.58 | 42.67 | 41.94 | 41.93 | 41.12 | 38.64 | 39.22 | 48.41 |
| PCC-S-C12 | 4.01 | PV-RCNN (STS-C) | 71.76 | 60.15 | 57.22 | 37.65 | 36.14 | 36.09 | 34.57 | 34.24 | 34.65 | 44.72 |
| RPKD-PV (CDTS) | 72.79 | 61.08 | 58.16 | 41.97 | 41.51 | 40.77 | 45.53 | 41.53 | 42.10 | 49.49 | ||
| PCC-S-C11 | 2.13 | PV-RCNN (STS-C) | 70.49 | 58.79 | 55.48 | 30.79 | 30.98 | 30.85 | 34.00 | 32.09 | 32.42 | 41.77 |
| RPKD-PV (CDTS) | 70.73 | 58.96 | 56.03 | 36.42 | 35.58 | 34.59 | 40.06 | 36.60 | 36.99 | 45.11 | ||
| PCC-S-C10 | 0.96 | PV-RCNN (STS-C) | 68.13 | 56.49 | 52.06 | 8.52 | 8.77 | 8.96 | 12.86 | 13.18 | 12.43 | 26.82 |
| RPKD-PV (CDTS) | 68.03 | 57.11 | 52.69 | 10.01 | 10.35 | 10.29 | 14.98 | 14.88 | 14.54 | 28.10 | ||